Tohban Report 2017-04-12

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 5 April 2017
End Date: 12 April 2017
Tohban: Juan Carlos Martinez Oliveros
Tohban email:
Next Tohban: Juan Camilo Buitrago Casas
List all reports


Solar Activity

Solar activity during was moderate at the start of the week and currently is at low levels. Currently there are two active regions on disk. Solar activity is expected to be very low with a slight chance for C-class activity through the forecast period (12-14 Apr).

How many GOES flares occurred?

 Flares above B, C, M, X class were     33     8     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were     13     7     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      8     3     0     0

There were RHESSI flares/GOES flares 54 / 41 over the time range 05-Apr-17 12-Apr-17

Memory Management

The SSR reached a maximum of ~70 percent during the week. Currently is about 60%

Spacecraft Status

During most of the week detectors 1, 3, 6 and 8 were active and taking data, but since 11-Apr-2017 only detectors 3 and 6 are active. The cold plate 1 and 2 temperatures are ~153.1K and 151.5K.

Decrease the attenuator threshold to 12.5% deadtime.

Currently the Cryocooler power is ~75.7 W and the acceleration is at ~11mg

Data Gaps

VC1 data gaps reported on 2017-04-05 and 2017-04-10

GAP START TIME              GAP END TIME                   GAP (SEC)
2017-04-06T01:05:00.000 -- 2017-04-06T05:40:00.000       16500.000
2017-04-06T06:50:00.000 -- 2017-04-07T00:00:00.000       61800.000
2017-04-07T00:00:00.000 -- 2017-04-08T00:00:00.000       86400.000
2017-04-08T00:00:00.000 -- 2017-04-09T00:00:00.000       86400.000
2017-04-09T00:00:00.000 -- 2017-04-10T00:00:00.000       86400.000
2017-04-10T00:00:00.000 -- 2017-04-11T00:00:00.000       86400.000
2017-04-11T00:00:00.000 -- 2017-04-12T00:00:00.000       86400.000

Detector issues

Detectors 1 and 8 were ramped to 0.0 V and their DIBs were powered off 11-Apr-2017 at 101-19:17 UTC.

Other notes

Spacecraft Management

Decimation Normal/Vigorous
HLAT Decimation Rear decimation
Night time data (fronts) plus/minus 4 minutes from daylight
Night time data (rears) plus/minus 4 minutes from daylight
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation logic enabled
Detector problems? See notes above.
Personal tools